Copper has been used for thousands of years, either as a pure metal or in alloys. It is frequently used today in the production of wires and cables. Copper can be obtained through smelting or recycling. Determine the energy associated with each of these processes in order to recycle 1.37 mol Cu. The smelting of copper occurs by the balanced chemical equation:
CuO(s) + CO(g) = Cu(s) +CO2(g)
where ΔH°f,CuO is = –155 kJ/mol. Assume the process of recycling copper is simplified to just the melting of the solid Cu starting at 25°C. The melting point of Cu is 1084.5°C with ΔH°fus = 13.0 kJ/mol and a molar heat capacity, cP,Cu = 24.5 J/mol·°C.

1. Enthalpy change for the reaction recovering Cu from CuO. answer in kj
2. Energy for recycling Cu. answer in kj
1
Expert's answer
2019-12-03T08:25:33-0500

Solution.

1.

Find the standard enthalpy of formation of Cu from CuO.

"\Delta H = (\Delta H (Cu) + \Delta H (CO2)) - (\Delta H (CuO) + \Delta H (CO))"

"\Delta H = (0 + (-393.51)) - (-155.00 + (-110.52)) = -127.99 \ \frac{kJ}{mol}"

Find the standard enthalpy of cu formation from CuO for 1.37 cu moles.

"\Delta H = 1.37 \times (-127.99) = - 175.35 \frac{kJ}{mol}"

2.

The energy for processing copper is found by the formula:

"\Delta H = \Delta H0 + \int \Delta Cp \Delta T"

"\Delta H = 13.0 + 24.5 \times (1084.5-25) = 38.96\frac{kJ}{mol}"

Answer:

"\Delta H = - 175.35 \frac{kJ}{mol}"

"\Delta H = 38.96 \frac{kJ}{mol}"